OpenStudy (anonymous):

SIMPLE QUESTION: Solve. Do checks. Explain. x + 4y = 2 -x + y = 8 Systems Of Elimination/Addition (I'm a visual learner)

OpenStudy (mathmate):

Add the equations to eliminate x. Solve for y. Back substitute to solve for x. Hint for check: x*y = -12

OpenStudy (texaschic101):

Here you go " visual learner " x + 4y = 2 -x + y = 8 -----------add 0 + 5y = 10 (as you can see, the x's canceled each other out) 5y = 10 (divide by 5 to get y by itself) 5y/5 = 10/5 y = 2 Now sub 2 in for y in either of the original equations x + 4y = 2 x + 4(2) = 2 x + 8 = 2 (subtract 8 from each side to get x by itself) x + 8 - 8 = 2 - 8 x = - 6 check.....you can use any of the original equations. Now just sub 2 in for y and -6 in for x. - x + y = 8 -(-6) + 2 = 8 6 + 2 = 8 8 = 8 (correct) ANSWER : x = -6 and y = 2